OpenStudy (anonymous):

If s = 1 over 4 unit and A = 80s2, what is the value of A, in square units? ____ square unit(s). (Input whole number only.) @dtan5457

OpenStudy (anonymous):

@Loser66

OpenStudy (dtan5457):

wording is tricking me up. not sure what it means

OpenStudy (loser66):

\(s=\dfrac{1}{4}\) and \(A = 80 s^2\). what is the value of A ? @dtan5457 try once

OpenStudy (dtan5457):

Is it 5? plug in 1/4 to a=80s^2 1/16 times 80=5 @Loser66

OpenStudy (loser66):

yup

OpenStudy (dtan5457):

thanks for the instructions @Loser66

OpenStudy (loser66):

not instructions, just interpret words problem

OpenStudy (dtan5457):

or giving me a chance to try it, appreciate it

OpenStudy (loser66):

:)